The difference of two numbers is 6. Their sum is 30. Find the two numbers.
Answer by checkley71(8403)

x-y=6 & x+y=30 or x=y+6 or y+6+y=30 or 2y=24 or y=12 & x-12=6 or x=18
proof 18-12=6 & 12+18=30